OpenStudy (anonymous):

Solve the equation sin (x° - 20°) = cos 42° for x, where 0 < x < 90. A. 62 B. 22 C. 68 D. 28

OpenStudy (campbell_st):

well cos(42) = sin(90 - 42) or cos(42) = sin(48) snow you can equation sin(x -20) = sin(48) and solve

OpenStudy (anonymous):

thanks